# Number of integral points on quadratic twists of elliptic curves

### v2

@@ Abstract.
We consider the integral points on the quadratic twists $E_{D}:y^{2}=x^{3}+D^{2}Ax+D^{3}B$ of the elliptic curve $E:y^{2}=x^{3}+Ax+B$ over $\mathbb{Q}$. For sufficiently large values of $D$, we prove that the number of integral points on $E_{D}$ admits the upper bound $\ll 4^{r}$, where $r$ denotes the Mordell-Weil rank of $E_{D}$.
@@ 1. Introduction
We will now restrict our attention to quadratic twist family of elliptic curves. For any squarefree integer $D$, consider the quadratic twist of $E$ by $D$,
@@
By [22], there exists an absolute constant $C$ such that
@@
where $r$ is the rank of $E_{D}/\mathbb{Q}$. The constant $C$ is computable, though is was not given in explicit form. Gross and Silverman [11] were the first to provide an explicit value of $C$, which is of order $10^{9}$. Subsequently, Chi, Lai, and Tan [5] refined this bound to $25$. Recently, Chan [4] further improved the bound to $3.8$ in the case of congruent number curves.
In this paper, we establish an improved bound $C=4$ for quadratic twist family of elliptic curves.
@@
Then for sufficiently large $\lvert D\rvert$ (depending on $A,B$),
$\left\lvert E_{D}(\mathbb{Z})\right\rvert\ll 4^{r},$
where $r$ is the rank of $E_{D}/\mathbb{Q}$.
@@
The bound $\ll$ is absolute, that is, it is independent of the elliptic curve $E$. We also mention the fact that the bound $\ll$ is effectively computable.
The principal tools utilized in the proof of Theorem 1.1 are gap principles and Diophantine approximation. We begin by partitioning $E_{D}(\mathbb{Z})$ into four subsets. We define the set of small points by
@@
The methods used to bound the four sets described above are all different.
To bound $E_{D}(\mathbb{Z})_{small}$, $E_{D}(\mathbb{Z})_{medium-small}$, and $E_{D}(\mathbb{Z})_{medium-large}$, we establish gap principles between integral points, and then apply bounds from spherical codes. The gap principle employed in our argument was originally formulated by Helfgott [12]. The precise form of the gap principle applied in this work is given in [1][Lemma 12].
To bound $E_{D}(\mathbb{Z})_{large}$, we once again employ the gap principle, but we will also employ the techniques from Diophantine approximation. Roughly speaking, we prove that large integral points yield rational approximations to a fixed algebraic number with exponent >2>2. We then invoke the quantitative Roth’s theorem to obtain a bound of large integral points. This kind of argument was originally given by [1], and subsequently refined by [4]. The form of Diophantine approximation applied in this work is analogous to [4][Lemma 6.1].
@@ 2. Notations
When we say ”for sufficiently large DD”, it means $D\geq D_{0}$ for some squarefree positive integer $D_{0}$ depending only on $E$. When we use the constants $c_{1},c_{2},\ldots$, they depend only on $E$.
When we write $f\ll g$, we mean $\lvert f\rvert\leq C\lvert g\rvert$, where $C$ is an absolute constant. Note that $C$ do not depend on $E$.
@@ 3. Preliminary lemmas
Apply Lemma 3.1 and [12][Lemma 4.1] for (1). Apply [12][Corollary 4.3] for (2). ∎
@@
Since $\phi_{D}$ is an isomorphism, $\hat{h}(P)=\hat{h}(\phi_{D}(P))$. By applying (2), we obtain (5). ∎
@@
By [12][Lemma 4.2], $E^{D}(\mathbb{Q})_{tors}\in\{0,\mathbb{Z}/2\mathbb{Z},(\mathbb{Z}/2\mathbb{Z})^{2}\}$ for sufficiently large $D$. Since $E^{D}$ and $E_{D}$ are isomorphic over $\mathbb{Q}$, the lemma is proved. ∎
@@ 4. Spherical codes
Let $r$ be a positive integer and $0<\theta<2\pi$ be an angle. Let $\Omega_{r}$ be the unit sphere in $\mathbb{R}^{r}$ and let $X$ be a finite subset of $\Omega_{r}$. We call $X$ a spherical code if $\langle x,y\rangle\leq\cos\theta$ for every $x,y\in X$. We write $A(r,\theta)$ for the maximum size of the spherical code $X$.
@@ 5. Bound for small points
First, we exclude torsion points from $E_{D}(\mathbb{Z})_{small}$. This is possible since by Lemma 3.5, $\lvert E_{D}(\mathbb{Q})_{tors}\rvert\leq 4$ for sufficiently large $D$.
@@
Fix $R\in E_{D}(\mathbb{Q})$. Suppose $P_{1},P_{2}\in\mathcal{S}(R)$ are distinct points and write $P_{1}-P_{2}=4Q$ for $Q\in E_{D}(\mathbb{Q})$. By Lemma 3.5, we may assume $Q$ is non-torsion for sufficiently large $D$. Then by Lemma 3.4,
$\hat{h}(P_{1}-P_{2})=16\hat{h}(Q)\geq 4\log D+16c_{3}\geq 3.5\log D$
@@
(8) $\cos\theta_{P_{1},P_{2}}=\frac{\hat{h}(P_{1})+\hat{h}(P_{1})-\hat{h}(P_{1}-P_{2})}{2\sqrt{\hat{h}(P_{1})\hat{h}(P_{2})}}\leq-\frac{0.5\log D}{3\log D}=-\frac{1}{6}<0.$
(8) shows that the angle between any two distinct points of $\mathcal{S}(R)$ is bounded from below by $\theta_{0}>\pi/2$. Therefore, Theorem 4.2 implies (7). ∎
@@ 8. Diophantine approximations
In this section, we prove Lemma 8.2. Roughly speaking, Lemma 8.2 says that large integral points give rational approximations to an algebraic number with exponent >2>2. By the celebrated Roth’s theorem [20], there are only finitely many such rational approximations, so we are able to bound large integral points.
@@
(14) $\frac{\log\lvert x(Q)-x(S)\rvert}{h(Q)}<-2.75.$
@@
$h(Q)\geq\frac{3(0.9-4\lambda-17\delta)}{4(10\lambda+\delta+0.1)}h(S)>5.77h(S).$
@@
By Lemma 3.8, Lemma 3.6, Lemma 3.7, and (23), we have
@@ 9. Quantitative Roth’s theorem
As mentioned above, Lemma 8.2 says that if $P=3Q+R$ is a very large integral point, then $x(Q)$ is a rational approximation to $x(S)$ with exponent >2.75>2.75. By Roth’s theorem, it is known that the number of such $x(Q)$ is finite. In order to bound the number, we need a quantitative version of Roth’s theorem.
@@ 10. Bound for large points
We first divide $E_{D}(\mathbb{Z})_{large}$ into cosets of $3E_{D}(\mathbb{Q})$. For any $R\in E_{D}(\mathbb{Q})$, define
@@
(32) $\lvert\mathcal{L}(R;S)^{**}\rvert\ll 1,\quad 3S=R.$
@@
$h(Q)\geq\max\{h(S),\log 2\},\quad\frac{\log\lvert x(S)-x(Q)\rvert}{h(Q)}<-2.75.$
Note that $x(S)$ is an algebraic number of degree at most 99. By Theorem 9.1 with $d\leq 9$ and $\epsilon=0.75$, (32) is proved. ∎

### v3

@@ Abstract.
We study integral points on the quadratic twists $E_{D}:y^{2}=x^{3}+D^{2}Ax+D^{3}B$ of a fixed elliptic curve $E:y^{2}=x^{3}+Ax+B$ over $\mathbb{Q}$. For sufficiently large squarefree positive integers $D$, we prove that the number of integral points on $E_{D}$ admits the upper bound $\ll 4^{r}$, where $r$ denotes the Mordell-Weil rank of $E_{D}$. The implied constant is absolute and effectively computable. The proof combines gap principles, bounds for spherical codes, and Diophantine approximation. As an application, we prove that the average number of integral points on the quadratic twist family is bounded.
Key words and phrases:
Elliptic curves, Quadrtic twists, Integral points
@@ 1. Introduction
We now restrict our attention to the quadratic twist family of a fixed elliptic curve. For each squarefree positive integer $D$, consider the quadratic twist of $E$ by $D$,
@@
Quadratic twist families are a natural setting for studying uniform bounds, as they depend on a single parameter $D$.
In this setting, one seeks bounds for $\lvert E_{D}(\mathbb{Z})\rvert$ that depend only on the Mordell-Weil rank of $E_{D}/\mathbb{Q}$. By [23], there exists an absolute constant $C$ such that
@@
where $r$ is the Mordell-Weil rank of $E_{D}/\mathbb{Q}$. The constant $C$ is computable, though is was not given in explicit form. Gross and Silverman [12] were the first to provide an explicit value of $C$, which is of order $10^{9}$. This was subsequently improved by Chi, Lai, and Tan [6], who refined the constant $C$ to about $25$. More recently, Chan [5] proved the sharper bound $C=3.8$ for the congruent number curve.
A principal motivation of this paper is to extend Chan’s result for the congruent number curve to general quadratic twist families. We prove that much of Chan’s method carries over to this setting, giving a bound with constant $C=4$. This improves the previous general bound of Chi, Lai, and Tan.
@@
Then there exists a constant $D(A,B)$, depending only on $A,B$, such that
$\left\lvert E_{D}(\mathbb{Z})\right\rvert\ll 4^{\mathrm{rank}(E_{D}/\mathbb{Q})}\quad\text{whenever}\quad D\geq D(A,B).$
@@
The bound $\ll$ is an absolute constant which is effectively computable.
Theorem 1.1 also has a natural arithmetic consequence. Recent work of Smith [25] shows that exponential moments of the Mordell-Weil rank are uniformly bounded in quadratic twist families. Specializing [25, Theorem 1.1] to our setting, we obtain the following theorem.
Theorem 1.3.
Let $E/\mathbb{Q}$ be an elliptic curve. Let $N(X)$ be the number of squarefree positive integers $D$ such that $0<D\leq X$. Then there exists a constant $C>0$ depending only on $E/\mathbb{Q}$ such that
$\frac{1}{N(X)}\sum_{\begin{subarray}{c}0<D\leq X\\ D\text{ squarefree}\end{subarray}}4^{\mathrm{rank}{(E_{D}/\mathbb{Q})}}\leq C,\quad X>0.$
Applying Theorem 1.1 and Theorem 1.3, we can prove that the average number of integral points on the quadratic twist family is bounded.
Corollary 1.4.
Let $E/\mathbb{Q}$ and $E_{D}/\mathbb{Q}$ be as in Theorem 1.1. Let $N(X)$ be the number of squarefree positive integers $D$ such that $0<D\leq X$. Then
$\limsup_{X\rightarrow\infty}\frac{1}{N(X)}\sum_{\begin{subarray}{c}0<D\leq X\\ D\text{ squarefree}\end{subarray}}\left\lvert E_{D}(\mathbb{Z})\right\rvert\ll C,$
where $C$ is a constant appearing in Theorem 1.3.
Proof.
Define
$S:=\sum_{\begin{subarray}{c}0<D\leq D(A,B)\\ D\text{ squarefree}\end{subarray}}\left\lvert E_{D}(\mathbb{Z})\right\rvert$.
Then
$\frac{1}{N(X)}\sum_{\begin{subarray}{c}0<D\leq X\\ D\text{ squarefree}\end{subarray}}\left\lvert E_{D}(\mathbb{Z})\right\rvert\ll\frac{S}{N(X)}+\frac{1}{N(X)}\sum_{\begin{subarray}{c}D(A,B)<D\leq X\\ D\text{ squarefree}\end{subarray}}4^{\mathrm{rank}{(E_{D}/\mathbb{Q})}}\leq\frac{S}{N(X)}+C.$
Letting $X\rightarrow\infty$ proves the theorem. ∎
Remark 1.5.
The bound $\ll$ in Corollary 1.4 is absolute. Consequently, if the constant $C$ in Theorem 1.3 could be chosen independently of $E/\mathbb{Q}$, then the same would hold for Corollary 1.4. In particular, this implies that the average number of integral points on the quadratic twist family is uniformly bounded.
Remark 1.6.
Recently, Browning and Chan [3] proved a much stronger average result. They proved that the average number of integral points on the quadratic twist family tends to 0, assuming the weak Hall-Lang conjecture if $x^{3}+Ax+B$ has precisely one root over $\mathbb{Q}$.
The proof of Theorem 1.1 combines gap principles, bounds for spherical codes, and Diophantine approximation, but these ingredients enter in different ways according to the height of the integral point. We begin by partitioning $E_{D}(\mathbb{Z})$ into four subsets:
@@
To bound $E_{D}(\mathbb{Z})_{small}$, $E_{D}(\mathbb{Z})_{medium-small}$, and $E_{D}(\mathbb{Z})_{medium-large}$, we establish gap principles between integral points, and then apply bounds from spherical codes. The form of the gap principle used in our argument was originally formulated by Helfgott [13], and subsequently refined by Alpoge [1].
To bound $E_{D}(\mathbb{Z})_{large}$, we once again employ the gap principle, but we will also employ the techniques from Diophantine approximation. Roughly speaking, we prove that large integral points yield rational approximations to a fixed algebraic number with exponent >2>2. We then invoke the quantitative Roth’s theorem to obtain a bound of large integral points. The Diophantine approximation of this kind originates from Alpoge [1], and subsequently refined by Chan [5].
@@ 2. Notations
When we say ”for sufficiently large DD”, it means $D\geq D(A,B)$ for some squarefree positive integer $D(A,B)$ depending only on $A,B$. When we use the constants $c_{1},c_{2},\ldots$, they depend only on $A,B$. When we write $f\ll g$, we mean $\lvert f\rvert\leq C\lvert g\rvert$, where $C$ is an absolute constant.
@@ 3. Preliminary lemmas
Apply Lemma 3.1 and [13, Lemma 4.1] for (1) and [13, Corollary 4.3] for (2). ∎
@@
Since $\phi_{D}$ is an isomorphism, $\hat{h}(P)=\hat{h}(\phi_{D}(P))$. Applying (2) gives (5). ∎
@@
By [13, Lemma 4.2], $E^{D}(\mathbb{Q})_{tors}\in\{0,\mathbb{Z}/2\mathbb{Z},(\mathbb{Z}/2\mathbb{Z})^{2}\}$ for sufficiently large $D$. Since $E^{D}$ and $E_{D}$ are isomorphic over $\mathbb{Q}$, the lemma is proved. ∎
@@ 4. Spherical codes
Let $r$ be a positive integer and $0<\theta<2\pi$ be an angle. Let $\Omega_{r}$ be the unit sphere in $\mathbb{R}^{r}$ and let $X$ be a finite subset of $\Omega_{r}$. We call $X$ a spherical code if $\langle x,y\rangle\leq\cos\theta$ for every distinct $x,y\in X$. We write $A(r,\theta)$ for the maximum size of the spherical code $X$.
@@ 5. Bound for small points
Let $\{P_{1},\ldots,P_{n}\}\subseteq\mathcal{S}(R)$ be maximal with the property that $P_{i}-P_{j}$ is non-torsion whenever $i\neq j$. For $i\neq j$, write $P_{i}-P_{j}=4S$ for some non-torsion point $S$. Then by Lemma 3.4,
$\hat{h}(P_{i}-P_{j})=16\hat{h}(S)\geq 4\log D+16c_{3}\geq 3.5\log D$
@@
(8) $\cos\theta_{P_{i},P_{j}}=\frac{\hat{h}(P_{i})+\hat{h}(P_{j})-\hat{h}(P_{i}-P_{j})}{2\sqrt{\hat{h}(P_{i})\hat{h}(P_{j})}}\leq-\frac{0.5\log D}{3\log D}=-\frac{1}{6}<0.$
(8) shows that the angle between any two distinct points of $\{P_{1},\ldots,P_{n}\}$ is bounded from below by $\theta_{0}>\pi/2$. By Theorem 4.2, $n$ is bounded by an absolute constant. By maximality, every $P\in\mathcal{S}(R)$ differs from some $P_{i}$ by a torsion point. By Lemma 3.5, $\lvert E_{D}(\mathbb{Q})_{tors}\rvert\leq 4$ for sufficiently large $D$, so we conclude that
$\lvert\mathcal{S}(R)\rvert\leq 4n\ll 1.$
@@ 8. Diophantine approximations
In this section, we prove Lemma 8.2. Roughly speaking, Lemma 8.2 says that large integral points give rational approximations to an algebraic number with exponent >2>2. By the celebrated Roth’s theorem [21], there are only finitely many such rational approximations, so we are able to bound large integral points. This kind of argument was originally given by Alpoge [1], and subsequently refined by Chan [5]. The form of Diophantine approximation applied in this work is analogous to [5][Lemma 6.1].
@@
(14) $\frac{\log\lvert x(Q)-x(S)\rvert}{h(Q)}<-2.7.$
@@
$h(Q)\geq\frac{3(0.9-4\lambda-17\delta)}{4(10\lambda+\delta+0.1)}h(S)>5.7h(S).$
@@
By Lemma 3.8, Lemma 3.6, Lemma 3.7, (23), and the relation PR=3QP-R=3Q, we have
@@ 9. Quantitative Roth’s theorem
As mentioned above, Lemma 8.2 says that if $P=3Q+R$ is a very large integral point, then $x(Q)$ is a rational approximation to $x(S)$ with exponent >2.7>2.7. By Roth’s theorem, it is known that the number of such $x(Q)$ is finite. In order to bound the number, we need a quantitative version of Roth’s theorem.
@@ 10. Bound for large points
We first divide $E_{D}(\mathbb{Z})_{large}$ into cosets of $3E_{D}(\mathbb{Q})$. For any point $R\in E_{D}(\mathbb{Q})$, define
@@
(32) $\lvert\mathcal{L}(R;S)^{**}\rvert\ll 1,\quad S\in\frac{1}{3}R.$
@@
$h(Q)\geq\max\{h(S),\log 2\},\quad\frac{\log\lvert x(S)-x(Q)\rvert}{h(Q)}<-2.7.$
Note that $x(S)$ is an algebraic number of degree at most 99. By Theorem 9.1 with $d\leq 9$ and $\epsilon=0.7$, (32) is proved. ∎